# SONAR: A Synthetic AI-Audio Detection Framework and Benchmark

### v3

@@ Abstract
Recent advances in Text-to-Speech (TTS) and Voice-Conversion (VC) using generative Artificial Intelligence (AI) technology have made it possible to generate high-quality and realistic human-like audio. This introduces significant challenges to distinguishing AI-synthesized speech from the authentic human voice and could raise potential issues of misuse for malicious purposes such as impersonation and fraud, spreading misinformation, deepfakes, and scams. However, existing detection techniques for AI-synthesized audio have not kept pace and often exhibit poor generalization across diverse datasets. In this paper, we introduce SONAR, a synthetic AI-Audio Detection Framework and Benchmark, aiming to provide a comprehensive evaluation for distinguishing cutting-edge AI-synthesized auditory content. SONAR includes a novel evaluation dataset sourced from 9 diverse audio synthesis platforms, including leading TTS providers and state-of-the-art TTS models. It is the first framework to uniformly benchmark AI-audio detection across both traditional and foundation model-based deepfake detection systems. Through extensive experiments, we reveal the generalization limitations of existing detection methods and demonstrate that foundation models exhibit stronger generalization capabilities, which can be attributed to their model size and the scale and quality of pretraining data. Additionally, we explore the effectiveness and efficiency of few-shot fine-tuning in improving generalization, highlighting its potential for tailored applications, such as personalized detection systems for specific entities or individuals. Code and dataset are available at https://github.com/Jessegator/SONAR
1 Introduction
@@
Recent advances in Text-to-Speech (TTS) and Voice-Conversion (VC) using Artificial Intelligence (AI) technology have made it possible to generate high-quality and realistic human-like audio efficiently [1, 2, 3, 4]. This introduces significant challenges in distinguishing AI-synthesized speech from the authentic human voice and could raise potential misuse for malicious purposes such as impersonation and fraud, spreading misinformation, and scams. For example, a deep fake AI voice of the US President Joe Biden was recently utilized in robocalls to advise them against voting [^1] https://www.cnn.com/2024/01/22/politics/fake-joe-biden-robocall/index.html, demonstrating how deepfakes can significantly manipulate public opinions and influence presidential elections. In response to such risks, the US Federal Communications Commission (FCC) now deems robot calls for election as illegal, which underscores the urgent need for enhanced detection of AI-synthesized audio.
@@
While TTS models are advancing rapidly, AI-synthesized audio detection techniques are not keeping pace. First, previous studies [5, 6] have highlighted the lack of generalization and robustness in these detection methods. Second, existing detection models [7, 6, 8, 9, 10] often take advantage of different audio features and evaluation datasets, complicating the comparison of their detection effectiveness. Third, a comprehensive evaluation to determine the effectiveness of these detection methods against the latest TTS models has not been conducted. This gap in research leaves a significant challenge in developing reliable detection techniques that can effectively counter the growing sophistication of AI-generated audio.
@@ 2 Related work
Text-to-Speech synthesis. Human voice synthesis is a significant challenge in the field of AI. State-of-the-art TTS synthesis approaches such as VALLE [4], AudioBox [1], VoiceBox [11], NaturalSpeech3 [4], and YourTTS [3] have demonstrated the possibility of generating high-quality, human-realistic audio with generative models trained on large datasets. Current TTS models can be classified into two primary categories: cascaded and end-to-end methods. Cascaded TTS models [12, 13, 14] typically employ a pipeline involving an acoustic model and a vocoder utilizing mel spectrograms as intermediary representations. To address the limitations associated with vocoders, end-to-end TTS models [15, 16] have been developed to jointly optimize both the acoustic model and vocoder. In practical applications, it is preferable to customize TTS systems to generate speech in any voice with limited accessible data. Consequently, there is increasing interest in zero-shot multi-speaker TTS techniques [17, 3, 2].
@@
AI-synthesized audio detection. Recent advancements in AI technology have significantly enhanced the ability to generate high-quality and realistic audio, calling for an urgent need for more robust and reliable detection methods. Several datasets have been developed to support research in this area. The ASVspoof challenges [18, 19, 20] are among the most notable, offering comprehensive datasets that cover a variety of attack vectors, including replay attacks, voice conversion, and directly synthesized audio. These resources aim to facilitate thorough evaluations of countermeasures against various spoofing techniques. In addition, newer datasets such as WaveFake [21] and LibriSeVoc [22] provide fake audio samples generated with state-of-the-art vocoders, offering diverse distributions to enhance the development of deepfake audio detection systems. By comparison, the In-the-Wild dataset [6] targets real-world applications by collecting deepfake audios from publicly accessible sources, capturing the complexity and diversity of manipulations encountered in everyday environments. Similarly, the SingFake dataset [6] focuses on the detection of synthetic singing voices, presenting unique challenges due to the musical content and variation in vocal expressions. These datasets are crucial for developing and testing next-generation AI-synthesized audio detection systems, pushing the boundaries of what is achievable in identifying and mitigating the threats posed by advanced audio synthesis technologies.
@@
Building upon these datasets, a significant body of research has focused on distinguishing AI-generated audio from genuine audio by designing advanced model architectures [8, 10, 7, 9] tailored for extracting different levels of representations of speech data for audio deepfake detection. Additionally, recent works [23, 24, 25] have leveraged speech foundation models for audio deepfake detection tasks. For instance, [23] and [24] fine-tune Wav2Vec2 [26] models on the ASVspoof dataset, while [25] uses Whisper as a front-end to extract audio features and trains various detection models based on these features, achieving state-of-the-art detection performance on the corresponding test datasets. However, none of these models have been evaluated on audio generated by the latest text-to-speech (TTS) models, leaving a gap in understanding their effectiveness against the most recent advancements in synthetic audio generation.
3 Evaluation dataset generalization and collection
@@
Figure 1: Overview of SONAR. Left: Audio deepfake data generation and collection. Right: Benchmark evaluation.
@@
Leveraging a set of diverse and high-quality speech data synthesis APIs and models, we create an evaluation dataset for synthetic AI-audio detection. Our approach incorporates two strategies: data generation and data collection. Our dataset includes AI-generated speech and audio from nine distinct sources. We perform speech data generation using one cutting-edge TTS service provider, OpenAI, and two open-sourced APIs, xTTS [27] and AudioGen [28]. For speech data collection, we leverage six state-of-the-art TTS models including Seed-TTS [29], VALL-E [4], PromptTTS2 [30], NaturalSpeech3 [31], VoiceBox [11], FlashSpeech [2]. Table 1 presents the details of our dataset generated by different audio generation models. We next detail our methods of generating and collecting these datasets.
@@
Table 1: Overview of our dataset with fake audios generated by various models. AudioGen lacks speaker and language information as it focuses on environmental sounds. The trainset sizes for OpenAI and Seed-TTS are unavailable due to the use of proprietary data. * denotes the samples that are directly collected from their demo page or provided test set due to the unavailability of their model checkpoints.
@@
Data generation. Our dataset generation involves OpenAI, xTTS, and AudioGen. Specifically, OpenAI currently provides voice choices from 6 different speakers. Using ChatGPT,we generate 100 different text prompts of varying lengths for each speaker, resulting in a total of 600 synthetic speech audios. xTTS supports synthetic speech generation given text prompts and reference speech. We select 6 speakers from the LibriTTS dataset [32] as the reference speech and also generate 600 text prompts with ChatGPT for each speaker, resulting in 600 synthetic speech audios. AudioGen can generate the corresponding environmental sound given a textual description of the acoustic scene. With AudioGen, we use ChatGPT to generate 100 text descriptions of the environment and background and obtain 100 AI-synthesized environmental sounds. Figure 1 (left) illustrates the data generation and collection process.
@@
To summarize, leveraging the details outlined above, we generate and collect a comprehensive evaluation dataset, encompassing a total of 2274 AI-synthesized audio samples produced by various TTS models. To the best of our knowledge, our dataset is by far the largest collection of fake audio generated by the latest TTS models. Note our motivation for collecting this dataset is for evaluation purposes. Additionally, we only include fake audio samples in this dataset since genuine audio samples can be easily collected from various sources (e.g., internet, self-recording, publicly available datasets, etc.). However, for convenience of evaluation, we also provide an equal number of real speech audio data sampled from the LibriTTS [32] clean-test set.
@@
4 Benchmarking AI-Audio Detection Models
@@ 4.1 Benchmarking setup
Model architectures. SONAR incorporates 11 models, including 5 state-of-the-art traditional audio deepfake detection models featuring various levels of input feature abstraction and 6 foundation models. Specifically, for the former, SONAR includes (1) AASIST [7], which processes raw waveform directly and utilizes graph neural networks and incorporates spectro-temporal attention mechanisms. (2) RawGAT-ST [9], which employs spectral and temporal sub-graphs along with a graph pooling strategy. (3) RawNet2 [8], which is a hybrid model combining CNN and GRU.(4) Spectrogram(Spec.)+ResNet [6], which transforms the audio to linear spectrogram using a 512-point Fast Fourier Transform (FFT) with a hop size of 10 ms. The spectrogram is then inputted into ResNet18 [33]. (5) LFCC-LCNN [10], which converts audio into Linear-Frequency Cepstral Coefficients (LFCC) for input into a CNN model. Specifically, 60-dimensional LFCCs are extracted from each utterance frame, with frame length set to 20ms and hop size 10ms. It extracts speech embedding directly from raw audio. These models collectively cover a broad spectrum of feature types and architectures, facilitating a detailed examination of their performance in deepfake audio detection applications. For foundation models, SONAR includes (1) Wave2Vec2 [26], which is pre-trained on 53k hours of unlabeled speech data. (2) Wave2Vec2BERT [34], which is pre-trained on 4.5M hours of unlabeled speech data covering more than 143 languages. (3) HuBERT [35], which is pretrained-on 60k hours of speech data. (4) CLAP [36], who is trained on a variety of audio-text pairs. (5) Whisper-small [37], and (6) Whisper-large [37]. Both Whispers are pre-trained on 680K hours of speech data covering 96 languages.
@@
Public datasets for training and testing. We consider three benchmark datasets for deepfake audio detection model training and testing as they are commonly used in the literature [38, 39]. Wavefake [21] collects deepfake audios from six vocoder architectures, including MelGAN [40], FullBand-MelGAN, MultiBand-MelGAN [41], HiFi-GAN [42], Parallel WaveGAN [43], and WaveGlow [44]. It consists of approximately 196 hours of generated audio files derived from the LJSPEECH [45] dataset. Similar to wavefake, LibriSeVoc [22] collects deepfake audios from six state-of-the-art neural vocoders including WaveNet[46], WaveRNN [47], Mel-GAN [41], Parallel WaveGAN [43], WaveGrad [48] and DiffWave [49] to generate speech samples derived from the widely used LibriTTS speech corpus [32], which is often utilized in text-to-speech research. Specifically, it consists of a total of 208.74 hours of synthesized samples. In-the-wild [5] comprises genuine and deepfake audio recordings of 58 politicians and other public figures gathered from publicly accessible sources, including social networks and video streaming platforms.
@@
Note that the test datasets are class-balanced, and the accuracy score is calculated using the EER threshold. Thus, we omit F1, precision, and recall scores from our evaluation results in the paper, though SONAR provides these metrics as well.
4.2 Results and analysis
4.2.1 How well can detection models generalize across datasets?
We first train all models on Wavefake training dataset and then evaluate the models on its own test set, LibriSeVoc test set, and In-the-wild dataset. Table 2 presents the evaluation results. Particularly, we make the following interesting observations.
Table 2: Generalization across existing audio deepfake datasets. All models are trained/finetuned on the Wavefake training set. Green and orange indicate the best and second-best performance, respectively.

### v4

# Where are we in audio deepfake detection? A systematic analysis over generative and detection models

@@ Abstract.
Recent advances in Text-to-Speech (TTS) and Voice-Conversion (VC) using generative Artificial Intelligence (AI) technology have made it possible to generate high-quality and realistic human-like audio. This poses growing challenges in distinguishing AI-synthesized speech from the genuine human voice and could raise concerns about misuse for impersonation, fraud, spreading misinformation, and scams. However, existing detection techniques for AI-synthesized audio have not kept pace and often fail to generalize across diverse datasets. In this paper, we introduce SONAR, a synthetic AI-Audio Detection Framework and Benchmark, aiming to provide a comprehensive evaluation for distinguishing cutting-edge AI-synthesized auditory content. SONAR includes a novel evaluation dataset sourced from 9 diverse audio synthesis platforms, including leading TTS providers and state-of-the-art TTS models. It is the first framework to uniformly benchmark AI-audio detection across both traditional and foundation model-based detection systems. Through extensive experiments, (1) we reveal the generalization limitations of existing detection methods and demonstrate that foundation models exhibit stronger generalization capabilities, which can be attributed to their model size and the scale and quality of pretraining data. (2) Speech foundation models demonstrate robust cross-lingual generalization capabilities, maintaining strong performance across diverse languages despite being fine-tuned solely on English speech data. This finding also suggests that the primary challenges in audio deepfake detection are more closely tied to the realism and quality of synthetic audio rather than language-specific characteristics. (3) We explore the effectiveness and efficiency of few-shot fine-tuning in improving generalization, highlighting its potential for tailored applications, such as personalized detection systems for specific entities or individuals. Code and dataset are available at https://github.com/Jessegator/SONAR.
Audio; Deepfake detection; Benchmark framework
1. Introduction
@@
Recent advances in Text-to-Speech (TTS) and Voice-Conversion (VC) using Artificial Intelligence (AI) technology have made it possible to generate high-quality and realistic human-like audio efficiently (Vyas et al., 2023; Ye et al., 2024; Casanova et al., 2022; Wang et al., 2023). This introduces significant challenges in distinguishing AI-synthesized speech from the authentic human voice and could raise potential misuse for malicious purposes such as impersonation and fraud, spreading misinformation, and scams. For example, a deepfake AI voice of the US President Joe Biden was recently utilized in robocalls to advise them against voting [^1] https://www.cnn.com/2024/01/22/politics/fake-joe-biden-robocall/index.html, demonstrating how deepfakes can significantly manipulate public opinions and influence presidential elections. In response to such risks, the US Federal Communications Commission (FCC) now deems robot calls for election as illegal, which underscores the urgent need for enhanced detection of AI-synthesized audio.
@@
While TTS models are advancing rapidly, AI-synthesized audio detection techniques are not keeping pace. First, previous studies (Müller et al., 2022; Zang et al., 2024) have highlighted the lack of generalization and robustness in these detection methods. Second, existing detection models (Jung et al., 2022; Zang et al., 2024; Tak et al., 2021b, a; Lavrentyeva et al., 2019) often take advantage of different audio features and evaluation datasets, complicating the comparison of their detection effectiveness. Third, a comprehensive evaluation to determine the effectiveness of these detection methods against the latest TTS models has not been conducted. This gap in research leaves a significant challenge in developing reliable detection techniques that can effectively counter the growing sophistication of AI-generated audio.
@@
  • Our evaluation of the generalization across languages suggests that speech foundation models demonstrate robust cross-lingual generalization capabilities, maintaining strong performance across diverse languages despite being fine-tuned solely on English speech data. This finding also suggests that the primary challenges in audio deepfake detection are more closely tied to the realism and quality of synthetic audio rather than language-specific characteristics.
@@ 2. Related work
Text-to-Speech synthesis. Human voice synthesis is a significant challenge in the field of AI. State-of-the-art TTS synthesis approaches such as VALLE (Wang et al., 2023), AudioBox (Vyas et al., 2023), VoiceBox (Le et al., 2024), NaturalSpeech3 (Wang et al., 2023), and YourTTS (Casanova et al., 2022) have demonstrated the possibility of generating high-quality, human-realistic audio with generative models trained on large datasets. Current TTS models can be classified into two primary categories: cascaded and end-to-end methods. Cascaded TTS models (Shen et al., 2018; Ren et al., 2019; Li et al., 2019) typically employ a pipeline involving an acoustic model and a vocoder utilizing mel spectrograms as intermediary representations. To address the limitations associated with vocoders, end-to-end TTS models (Kim et al., 2021; Liu et al., 2022) have been developed to jointly optimize both the acoustic model and vocoder. In practical applications, it is preferable to customize TTS systems to generate speech in any voice with limited accessible data. Consequently, there is increasing interest in zero-shot multi-speaker TTS techniques (Cooper et al., 2020; Casanova et al., 2022; Ye et al., 2024).
@@
AI-synthesized audio detection. Recent advancements in AI technology have significantly enhanced the ability to generate high-quality and realistic audio, creating an urgent need for more robust and reliable detection methods. Several comprehensive datasets have been developed to support research in this field. The ASVspoof challenges (Wu et al., 2017; Wang et al., 2020; Nautsch et al., 2021; Todisco et al., 2019; Liu et al., 2023) are among the most notable, offering datasets that cover various attack vectors including replay attacks, voice conversion, and directly synthesized audio. Building upon this foundation, the ADD challenges (Yi et al., 2022, 2023) and CFAD dataset (Ma et al., 2024) introduce more challenging scenarios by incorporating background noise, disturbances, and partial audio manipulations, though both are limited to Chinese language content. To address specific technical challenges, WaveFake (Frank and Schönherr, 2021) and LibriSeVoc (Sun et al., 2023) focus on state-of-the-art vocoder-generated samples, while Codecfake (Xie et al., 2024) specifically targets the detection of audio generated by Large Language Model (LLM)-based systems that operate directly with discrete neural codecs. The MLAAD dataset (Müller et al., 2024) provides broad linguistic coverage, incorporating fake audio from 82 Text-to-Speech models across 38 languages. For real-world applications, the In-the-Wild dataset (Müller et al., 2022) collects deepfake audio from publicly accessible sources, capturing the complexity of manipulations encountered in everyday environments. Similarly, the SingFake dataset (Zang et al., 2024) addresses the unique challenges of synthetic singing voice detection, where musical content and varied vocal expressions add additional complexity. Collectively, these datasets are crucial for developing and evaluating next-generation AI-synthesized audio detection systems, advancing our ability to identify and mitigate threats posed by sophisticated audio synthesis technologies.
@@
Building upon these datasets, a significant body of research has focused on distinguishing AI-generated audio from genuine audio by designing advanced model architectures (Tak et al., 2021b; Lavrentyeva et al., 2019; Jung et al., 2022; Tak et al., 2021a) tailored for extracting different levels of representations of speech data for audio deepfake detection. Additionally, recent works (Wang and Yamagishi, 2021; Tak et al., 2022; Kawa et al., 2023) have leveraged speech foundation models for audio deepfake detection tasks. For instance, (Wang and Yamagishi, 2021) and (Tak et al., 2022) fine-tune Wav2Vec2 (Baevski et al., 2020) models on the ASVspoof dataset, while (Kawa et al., 2023) uses Whisper as a front-end to extract audio features and trains various detection models based on these features, achieving state-of-the-art detection performance on the corresponding test datasets. However, none of these models have been evaluated on audio generated by the latest TTS models, leaving a gap in understanding their effectiveness against the most recent advancements in synthetic audio generation.
SONAR offers several key advantages: (1) it provides comprehensive empirical evaluation of both traditional detection models and foundation models, including those not previously explored; (2) it incorporates a diverse range of audio sources, from web demos to compliant commercial APIs, while ensuring adherence to usage policies; (3) it examines the impact of model architecture, training data, and few-shot fine-tuning on detection performance; and (4) it reveals critical insights about the trade-off between fine-tuning effectiveness and model generalization. These contributions make SONAR a more extensive and systematic benchmark for audio deepfake detection.
3. Evaluation dataset generalization and collection
@@
Figure 1. Overview of SONAR. Left: Audio deepfake data generation and collection. Right: Benchmark evaluation.
@@
Leveraging a set of diverse and high-quality speech data synthesis APIs and models, we create an evaluation dataset for synthetic AI-audio detection. Our approach incorporates two strategies: data generation and data collection. Our dataset includes AI-generated speech and audio from nine distinct sources. We perform speech data generation using one cutting-edge TTS service provider, OpenAI, and two open-sourced APIs, xTTS (Casanova et al., 2024) and AudioGen (Kreuk et al., 2022).
For speech data collection, we leverage six state-of-the-art TTS models including Seed-TTS (Anastassiou et al., 2024), VALL-E (Wang et al., 2023), PromptTTS2 (Leng et al., 2023), NaturalSpeech3 (Ju et al., 2024), VoiceBox (Le et al., 2024), FlashSpeech (Ye et al., 2024). Table 1 presents the details of our dataset generated by different audio generation models. We next detail our methods of generating and collecting these datasets.
@@
Table 1. Overview of our dataset with fake audios generated by various models. AudioGen lacks speaker and language information as it focuses on environmental sounds. The trainset sizes for OpenAI and Seed-TTS are unavailable due to the use of proprietary data. * denotes the samples that are directly collected from their demo page or provided test set due to the unavailability of their model checkpoints.
@@
Data generation. Our dataset generation involves OpenAI, xTTS, and AudioGen. Specifically, OpenAI currently provides voice choices from 6 different speakers. Using ChatGPT, we generate 100 different text prompts of varying lengths for each speaker, resulting in a total of 600 synthetic speech audios. xTTS supports synthetic speech generation given text prompts and reference speech. We select 6 speakers from the LibriTTS dataset (Zen et al., 2019) as the reference speech and also generate 100 text prompts with ChatGPT for each speaker, resulting in 600 synthetic speech audios. To evaluate whether speech detection models can generalize to AI-synthesized environmental sound, we also include a subset consisting of AI-synthesized environmental sound. AudioGen can generate the corresponding environmental sound given a textual description of the acoustic scene. With AudioGen, we use ChatGPT to generate 100 text descriptions of the environment and background and obtain 100 AI-synthesized environmental sounds. Figure 1 (left) illustrates the data generation and collection process.
@@
To summarize, leveraging the outlined details, we generate and collect a comprehensive evaluation dataset, encompassing a total of 2274 AI-synthesized audio samples produced by various TTS models. To the best of our knowledge, our dataset is by far the largest collection of fake audio generated by the latest TTS models. Note our motivation for collecting this dataset is for evaluation purposes. Additionally, we only include fake audio samples in this dataset since genuine audio samples can be easily collected from various sources (e.g., internet, self-recording, publicly available datasets, etc.). However, for convenience of evaluation, we also provide an equal number of real speech audio data sampled from the LibriTTS (Zen et al., 2019) clean-test set.
@@
4. Benchmarking AI-Audio Detection Models
@@ 4.1. Benchmarking setup
Model architectures. SONAR incorporates 11 models, including 5 state-of-the-art traditional audio deepfake detection models featuring various levels of input feature abstraction and 6 foundation models. Specifically, for the former, SONAR includes (1) AASIST (Jung et al., 2022), which processes raw waveform directly and utilizes graph neural networks and incorporates spectro-temporal attention mechanisms. (2) RawGAT-ST (Tak et al., 2021a), which employs spectral and temporal sub-graphs along with a graph pooling strategy. (3) RawNet2 (Tak et al., 2021b), which is a hybrid model combining CNN and GRU.(4) Spectrogram(Spec.)+ResNet (Zang et al., 2024), which transforms the audio to linear spectrogram using a 512-point Fast Fourier Transform (FFT) with a hop size of 10 ms. The spectrogram is then inputted into ResNet18 (He et al., 2016). (5) LFCC-LCNN (Lavrentyeva et al., 2019), which converts audio into Linear-Frequency Cepstral Coefficients (LFCC) for input into a CNN model. Specifically, 60-dimensional LFCCs are extracted from each utterance frame, with frame length set to 20ms and hop size 10ms. It extracts speech embedding directly from raw audio. These models collectively cover a broad spectrum of feature types and architectures, facilitating a detailed examination of their performance in deepfake audio detection applications. For foundation models, SONAR includes (1) Wave2Vec2 (Baevski et al., 2020), which is pre-trained on 53k hours of unlabeled speech data. (2) Wave2Vec2BERT (Barrault et al., 2023), which is pre-trained on 4.5M hours of unlabeled speech data covering more than 143 languages. (3) HuBERT (Hsu et al., 2021), which is pretrained-on 60k hours of speech data. (4) CLAP (Elizalde et al., 2023), which is trained on a variety of audio-text pairs. (5) Whisper-small (Radford et al., 2023), and (6) Whisper-large (Radford et al., 2023). Both Whispers are pre-trained on 680K hours of speech data covering 96 languages.
@@
Public datasets for training and testing. We consider five benchmark datasets for deepfake audio detection model training and testing as they are commonly used in the literature (Kawa et al., 2022b, a). Wavefake (Frank and Schönherr, 2021) collects deepfake audios from six vocoder architectures, including MelGAN (Kumar et al., 2019), FullBand-MelGAN, MultiBand-MelGAN (Yang et al., 2021), HiFi-GAN (Kong et al., 2020a), Parallel WaveGAN (Yamamoto et al., 2020), and WaveGlow (Prenger et al., 2019). It consists of approximately 196 hours of generated audio files derived from the LJSPEECH (Ito and Johnson, 2017) dataset. Similar to wavefake, LibriSeVoc (Sun et al., 2023) collects deepfake audios from six state-of-the-art neural vocoders including WaveNet(Van Den Oord et al., 2016), WaveRNN (Kalchbrenner et al., 2018), Mel-GAN (Yang et al., 2021), Parallel WaveGAN (Yamamoto et al., 2020), WaveGrad (Chen et al., 2020) and DiffWave (Kong et al., 2020b) to generate speech samples derived from the widely used LibriTTS speech corpus (Zen et al., 2019), which is often utilized in text-to-speech research. Specifically, it consists of a total of 208.74 hours of synthesized samples. In-the-wild (Müller et al., 2022) comprises genuine and deepfake audio recordings of 58 politicians and other public figures gathered from publicly accessible sources, including social networks and video streaming platforms. MLAAD (Müller et al., 2024) consists of fake audios created using 82 TTS models, covering 38 different languages. We use this dataset to evaluate the cross-lingual generalization of different detection models. To further evaluate generalization capabilities on voice-converted audio, we utilize a subset of the ASVSpoof2019 development set (Wang et al., 2020), which contains synthetic speech generated through voice-conversion techniques. All input audios are resampled to a 16kHz sampling rate and converted into raw waveforms consisting of 64,000 samples (approximately 4 seconds). Audios longer than 4 seconds are randomly trimmed, while those shorter than 4 seconds are repeated and padded to meet the 4-second duration.
@@
Note that the test datasets are class-balanced, and the accuracy score is calculated using the EER threshold. Thus, we omit F1, precision, and recall scores from our evaluation results in the paper.
Table 2. Hyperparemeters
config value
optimizer Adam
optimizer momentum $\beta_{1}=0.9$, $\beta_{2}=0.999$
weight decay 1e-4
epochs 40
warmup epochs 0
scheduler cosine decay
Implementation Details. Table 2 presents the hyperparameters for training AASIST, RawNet2, RawGAT-ST, LCNN, and Spec.+ResNet. We train AASIST, RawNet2, and RawGAT-ST with a learning rate of 0.0001 and LCNN and Spec.+ResNet with a learning rate of 0.0003. The batch size for AASIST, RawNet2, RawGAT-ST, LCNN, and Spec.+ResNet are 64, 256, 32, 512, and 256, respectively. All input audios are resampled to a 16kHz sampling rate and converted into raw waveforms consisting of 64,000 samples (approximately 4 seconds). Audios longer than 4 seconds are randomly trimmed, while those shorter than 4 seconds are repeated and padded to meet the 4-second duration.
4.2. Results and analysis
4.2.1. How well can detection models generalize across datasets?
We first train all models on Wavefake training dataset and then evaluate the models on its own test set, LibriSeVoc test set, and In-the-wild testset. Table 3 presents the evaluation results. Particularly, we make the following interesting observations.
Table 3. Generalization across existing audio deepfake datasets. All models are trained/finetuned on the Wavefake training set. Green and orange indicate the best and second-best performance, respectively.